# Inversion sequences growing on the left

### v1

@@ Abstract
This work concerns a construction of pattern-avoiding inversion sequences from right to left, called a generating tree growing on the left. We first apply this construction to inversion sequences avoiding 201 and 210, resulting in a new way of computing their generating function. We then use a slightly modified construction to compute the generating function of inversion sequences avoiding 010 and 102, which was only conjectured before.
[…]
biblio.bib
@@
Benjamin Testart [^1] Université de Lorraine, CNRS, Inria, LORIA, F 54000 Nancy, France
@@ 1.1 Basic definitions
For all $n\geqslant 0$, let $\mathcal{I}_{n}$ be the set of inversion sequences [Mansour_Shattuck_2015, Corteel_Martinez_Savage_Weselcouch_2016] of size $n$, that is the set of sequences $\sigma=(\sigma_{1},\dots,\sigma_{n})$ such that $\sigma_{i}\in\{0,\dots,i-1\}$ for all $i\in\{1,\dots,n\}$. Let $\mathcal{I}=\coprod_{n\geqslant 0}\mathcal{I}_{n}$ be the set of all inversion sequences.
@@
We say that an inversion sequence $\sigma\in\mathcal{I}_{n}$ contains a pattern $\rho$ of length $k\leqslant n$ if there exists a subsequence of $k$ entries of $\sigma$ whose values appear in the same order as the values of $\rho$. Such a subsequence is called an occurrence of $\rho$. An inversion sequence avoids the pattern $\rho$ if it does not contain $\rho$. For instance, the sequence $(0,1,0,3,0,2,3)$ avoids the pattern 110, but it contains the pattern 021 since the subsequence $(0,3,2)$ is an occurrence of 021. If $P$ is a set of patterns, we denote by $\mathcal{I}(P)$ the set of inversion sequences avoiding all patterns in $P$, and we denote by $\mathcal{I}_{n}(P)$ its subset of sequences of size $n$.
@@ 1.2 Context and summary of results
Pattern-avoiding inversion sequences were introduced in [Mansour_Shattuck_2015] and [Corteel_Martinez_Savage_Weselcouch_2016], by analogy with pattern-avoiding permutations. Afterwards, many more articles studied inversion sequences avoiding patterns of length three, see [Martinez_Savage_2018, Beaton_Bouvel_Guerrini_Rinaldi_2019, Yan_Lin_2020, Kotsireas_Mansour_Yildirim_2024, Chen_Lin, Pantone201+210, Testart2024], among others.
@@
Our work is concerned with two distinct classes of pattern-avoiding inversion sequences: $\mathcal{I}(201,210)$ and $\mathcal{I}(010,102)$. The class $\mathcal{I}(201,210)$ is counted by sequence A212198 in [OEIS]. It has an algebraic generating function, which was found in both [Chen_Lin] and [Pantone201+210], using two different constructions. Chen and Lin [Chen_Lin] use a decomposition around the rightmost saturated [^2] In an inversion sequence $\sigma$, the entry $\sigma_{i}$ is said to be ”saturated” if $\sigma_{i}=i-1$ (since $i-1$ is the largest value which could appear at position $i$). entry of inversion sequences, with one parameter counting the number of saturated entries. The construction used by Pantone [Pantone201+210] is a generating tree in one integer label and two boolean labels, and it involves some ”commitments” to future decisions.
@@
The class $\mathcal{I}(010,102)$ is counted by sequence A374553 in [OEIS]. This sequence was computed in polynomial time in [Testart2024] using a decomposition involving $\{010,102\}$-avoiding words, and it was conjectured to have an algebraic generating function in [Pantone201+210] using the software [guessfunc].
@@
In Section 2, we describe a construction for inversion sequences, which we call the ”generating tree growing on the left”. This construction first appeared in [Corteel_Martinez_Savage_Weselcouch_2016] to enumerate the classes $\mathcal{I}(101)$ and $\mathcal{I}(110)$. A modified construction based on this one also appears in [Lin_Fu_2021] to enumerate inversion sequences avoiding the vincular pattern $\underline{12}0$. By considering the generating tree growing on the left for inversion sequences avoiding 201 and 210, we find a new and simpler way of computing the generating function of $\mathcal{I}(201,210)$, in Section 3. We then modify this construction to apply it to inversion sequences avoiding 010 and 102 in Section 4. This allows us to compute the generating function of $\mathcal{I}(010,102)$ which was conjectured in [Pantone201+210].
@@
As it is often the case with generating trees, we first turn the succession rule describing the tree into a functional equation for a multivariate generating function; we then solve this equation using the kernel method [GFforGT]. The equation we obtain for $\mathcal{I}(010,102)$ contains two evaluations $A(t,1)$ and $A(t,2)$ of a bivariate series $A$. This occurs because the succession rule produces labels with multiplicities that are powers of 2. In the case of $\mathcal{I}(201,210)$, the functional equation involves two catalytic variables, but it is easily solved by simply applying the kernel method twice.
@@ 1.3 Notation
For any integers $a,b\in\mathbb{Z}$, we denote the integer interval $[a,b]=\{k\in\mathbb{Z}\;:\;a\leqslant k\leqslant b\}$. We denote by $\varepsilon$ the empty sequence, that is the only sequence of size 00. We define $\chi$ to be the characteristic function of a statement, that is $\chi(\text{S})=\begin{cases}0&\text{if \; S is false},\\ 1&\text{if \; S is true}.\end{cases}$
@@ 2 Generating trees for inversion sequences growing on the left
A generating tree [West1, West2] is a rooted, labelled tree such that the label of each node determines its number of children, and their labels. A generating tree can be defined by a succession rule
$\Omega=\left\{\begin{array}[]{rclll}a\\ \ell&\leadsto&i^{m(\ell,i)}&\text{for}&i\in C(\ell)\\ \end{array}\right.$
@@
Generating trees are used to solve enumeration problems by identifying combinatorial objects with the nodes of a tree, so that each object of size $n$ corresponds to a node at level $n$ in the tree. This is useful when many nodes are roots of isomorphic subtrees, as the combinatorial objects corresponding to these nodes may all be given the same label. A succession rule for the generating tree then induces a recurrence relation on the number of objects of each size (which usually involves some parameters corresponding to the labels of the tree), or equivalently a functional equation for the (multivariate) generating function. For a more thorough introduction to generating trees in the context of pattern-avoiding inversion sequences, see [Testart2024, Section 2.1].
@@
$\Omega_{\text{Left}}=\left\{\begin{array}[]{rclll}(0)\\ (k)&\leadsto&(i+1)^{\binom{k}{i}}&\text{for}&i\in[0,k]\\ \end{array}\right.$
@@
The simplest construction of inversion sequences is the generating tree growing sequences on the right, used in e.g. [Kotsireas_Mansour_Yildirim_2024] or [Testart2024, Section 2]. In comparison, growing sequences on the left is not the most practical way to construct the set of inversion sequences (it is not even obvious that the tree described by $\Omega_{\text{Left}}$ has $n!$ nodes at each level $n$). However, this construction becomes simpler for some classes of pattern-avoiding inversion sequences. For instance, [Corteel_Martinez_Savage_Weselcouch_2016] proves that the generating trees growing on the left for the classes $\mathcal{I}(101)$ and $\mathcal{I}(110)$ are both described by the succession rule
$\Omega_{\text{pCat}}=\left\{\begin{array}[]{rclll}(0)\\ (k)&\leadsto&(k+1)\\ &&(i)^{i}&\text{for}&i\in[1,k].\\ \end{array}\right.$
The generating tree growing on the left for inversion sequences lacks a property that is typically desired when working with objects that avoid patterns: it does not automatically restrict to a generating tree for families of pattern-avoiding inversion sequences. Specifically, restricting the tree to keep only sequences which avoid some patterns may disconnect the tree. This occurs because an inversion sequence which contains a pattern $\rho$ may have children which avoid $\rho$. For instance, the sequence $(0,0,2,1)$ avoids the pattern 010, but its parent sequence is $(0,1,0)$. However, we show in Proposition 1 below that this problem only occurs with patterns which contain several zeros. Even then, it can be worked around in some cases. For instance, we slightly modify our construction in Section 4 to describe a generating tree for the class $\mathcal{I}(010,102)$.
@@
Let $\rho$ be a pattern of length $k\geqslant 1$ which contains a single zero, let $\sigma$ be an inversion sequence which contains an occurrence $(\sigma_{q_{i}})_{i\in[1,k]}$ of $\rho$, and let $\tau$ be a child of $\sigma$ in the generating tree growing on the left. Then $\tau$ contains the pattern $\rho$.
@@
Since $\rho$ contains a single zero, $(\sigma_{q_{i}})_{i\in[1,k]}$ contains at most one zero. It follows that $(\sigma_{q_{i}})_{i\in[1,k]}$ is order-isomorphic to $(\tau_{q_{i}+1})_{i\in[1,k]}$ (every nonzero value is increased by 1, and the hypothetical zero can either increase by 1 or remain a zero), hence $\tau$ contains the pattern $\rho$. ∎
@@ 3 Inversion sequences avoiding 201 and 210
For all $n\geqslant 0$ and $\sigma\in\mathcal{I}_{n}(201,210)$, let $L(\sigma)=\{i\in[1,n]\,:\,\forall j\leqslant i,\;\sigma_{j}=0\}$ be the set of leading zeros of $\sigma$, and let $R(\sigma)=\mathbf{Zeros}(\sigma)\backslash L(\sigma)$. Let $\ell(\sigma)=|L(\sigma)|$ be the number of leading zeros of $\sigma$. If $\ell(\sigma)<n$ (i.e. if $\sigma$ is not constant), then $\sigma_{\ell(\sigma)+1}$ is the value of the first nonzero entry of $\sigma$. Notice that $L(\sigma)$ (resp. $R(\sigma)$) is the set of zeros of $\sigma$ to the left (resp. right) of position $\ell(\sigma)+1$.
@@
The proposition above establishes that in the generating tree growing on the left for $\mathcal{I}(201,210)$, the children of $\sigma$ are the sequences such that the subset $Z$ of zeros to which the value 1 is added must contain either all or none of $R(\sigma)$. Any subset of $L(\sigma)$ may intersect with $Z$ without creating an occurrence of 201 or 210. It follows that the number of children of $\sigma$ is $2^{\ell(\sigma)+\chi(R(\sigma)\neq\emptyset)}$.
@@
    • If $Z=[i,\ell(\sigma)]$ for some $i\in[1,\ell(\sigma)+1]$, then $L(\tau)=[1,i]$ and $R(\tau)=\emptyset$, therefore $s(\tau)=i$.
    • Otherwise $R(\tau)\neq\emptyset$. We have $s(\tau)=i+1$, where $i=\ell(\tau)=\mathbf{min}(Z)\in[1,\ell(\sigma)-1]$. There are $12^{s(\sigma)-i}-$ 1 possible sets $Z$ in that case (that is all sets of the form $Z=\{i\}\sqcup Z^{\prime}$ for some $Z^{\prime}\subsetneq[i+1,\ell(\sigma)]$).
@@
    • Otherwise, $Z$ is not empty nor consists of a set of final zeros of $\sigma$. This implies that $R(\tau)\neq\emptyset$, and that $s(\tau)=i+1$, where $i=\ell(\tau)=\mathbf{min}(Z)\in[1,\ell(\sigma)]$. There are $12^{s(\sigma)-i}-$ 1 possible sets $Z$ in that case: either $Z=\{i\}\sqcup Z^{\prime}$ for some $Z^{\prime}\subseteq[i+1,\ell(\sigma)]$, or $Z=\{i\}\sqcup Z^{\prime}\sqcup R(\sigma)$ for some $Z^{\prime}\subsetneq[i+1,\ell(\sigma)]$). ∎
@@
We apply the kernel method. The kernel in (1) is $1+\frac{tu^{2}}{2-u}$. To cancel the kernel, we substitute $u$ with $U(t)=\frac{1-\sqrt{1-8t}}{2t}$. This yields the equation

### v2

@@ Abstract
This work concerns a construction of pattern-avoiding inversion sequences from right to left we call the generating tree growing on the left. We first apply this construction to inversion sequences avoiding 201 and 210, resulting in a new way of computing their generating function. We then use a slightly modified construction to compute the generating function of inversion sequences avoiding 010 and 102, which was only conjectured before. These generating functions are algebraic in both instances. We end by discussing how the generating tree growing on the left can be applied in a more general setting.
@@
Benjamin Testart [^1] Université de Lorraine, CNRS, Inria, LORIA, F-54000 Nancy, France
@@ 1.1 Basic definitions
For all $n\geqslant 0$, let $\mathcal{I}_{n}$ be the set of inversion sequences [MS15, CMSW16] of size $n$, that is, the set of sequences $\sigma=(\sigma_{1},\dots,\sigma_{n})$ such that $\sigma_{i}\in\{0,\dots,i-1\}$ for all $i\in\{1,\dots,n\}$. Let $\mathcal{I}=\coprod_{n\geqslant 0}\mathcal{I}_{n}$ be the set of all inversion sequences.
@@
We say that an inversion sequence $\sigma\in\mathcal{I}_{n}$ contains a pattern $\rho$ of length $k\leqslant n$ if there exists a subsequence of $k$ entries of $\sigma$ whose values appear in the same order as the values of $\rho$. Such a subsequence is called an occurrence of $\rho$. An inversion sequence avoids the pattern $\rho$ if it does not contain $\rho$. For instance, the sequence $(0,1,0,3,0,2,3)$ avoids the pattern 110, but it contains the pattern 021 since the subsequence $(0,3,2)$ is an occurrence of 021. If $P$ is a set of patterns, we denote by $\mathcal{I}(P)$ the set of inversion sequences avoiding all patterns in $P$, and by $\mathcal{I}_{n}(P)$ its subset of sequences of size $n$.
@@ 1.2 Context and summary of results
Pattern-avoiding inversion sequences were introduced in [MS15] and [CMSW16], by analogy with pattern-avoiding permutations. Afterwards, many more articles studied inversion sequences avoiding patterns of length three, see [MS18, BBGR19, YL20, KMY24, CL24, Pan24, Tes24], among others.
@@
Our work is concerned with two distinct classes of pattern-avoiding inversion sequences: $\mathcal{I}(201,210)$ and $\mathcal{I}(010,102)$. The class $\mathcal{I}(201,210)$ is counted by sequence A212198 in [OEIS24]. It has an algebraic generating function, which was found in both [CL24] and [Pan24], using two different constructions. Chen and Lin [CL24] use a decomposition around the rightmost saturated [^2] In an inversion sequence $\sigma$, the entry $\sigma_{i}$ is said to be “saturated” if $\sigma_{i}=i-1$ (since i1i-1 is the largest value which could appear at position $i$). entry of inversion sequences, with one parameter counting the number of saturated entries. The construction used by Pantone [Pan24] is a generating tree in one integer parameter and two boolean parameters, and it involves some “commitments” to future decisions.
@@
The class $\mathcal{I}(010,102)$ is counted by sequence A374553 in [OEIS24]. This sequence was computed in polynomial time in [Tes24] using a decomposition involving $\{010,102\}$-avoiding words, and it was conjectured to have an algebraic generating function in [Pan24] using the software [Pan21].
@@
In Section 2, we describe a construction for inversion sequences, which we call the “generating tree growing on the left”. This construction first appeared in [CMSW16] to enumerate the classes $\mathcal{I}(101)$ and $\mathcal{I}(110)$. A modified construction based on this one also appears in [LF21] to enumerate inversion sequences avoiding the vincular pattern $\underline{12}0$.
@@
As it is often the case with generating trees, our general approach in this paper is to first look for a simple succession rule describing the tree (up to isomorphism), then turn this rule into a functional equation for a multivariate generating function, then solve this equation with the kernel method [BBDFGG02].
In Section 3, we consider the generating tree growing on the left for inversion sequences avoiding 201 and 210. Our succession rule corresponds to a functional equation which contains two evaluations $A(t,1)$ and $A(t,2)$ of a bivariate series $A$. This is quite unusual, and it occurs because the succession rule produces labels with multiplicities that are powers of 2. Regardless, this equation can still be solved with the kernel method, resulting in a new and simpler way of computing the generating function of $\mathcal{I}(201,210)$.
In Section 4, we modify the generating tree growing on the left construction to apply it to inversion sequences avoiding 010 and 102. In this case, we obtain a functional equation involving two catalytic variables, which can be solved by simply applying the kernel method twice. This allows us to compute the generating function of $\mathcal{I}(010,102)$, confirming a conjecture of [Pan24].
In Section 5, we briefly explore how widely the generating tree growing on the left could be applied to other similar problems. To that end, we present some succession rules for inversion sequences avoiding any pattern of length 3.
@@ 1.3 Notation
[…]
@@ 2 Generating trees for inversion sequences growing on the left
A generating tree [Wes95, Wes96] is a rooted, labelled tree such that the label of each node determines its number of children, and their labels. A generating tree can be defined by a succession rule
[…]
@@
Generating trees are used to solve enumeration problems by identifying combinatorial objects with the nodes of a tree, so that each object of size $n$ corresponds to a node at level $n$ in the tree. This is useful when many nodes are roots of isomorphic subtrees, as the combinatorial objects corresponding to these nodes may all be given the same label. A succession rule for a generating tree then induces a recurrence relation on the number of objects of each size (which usually involves some parameters corresponding to the labels of the tree), or equivalently a functional equation for a (multivariate) generating function. For a more thorough introduction to generating trees in the context of pattern-avoiding inversion sequences, see [Tes24, Section 2.1].
@@
[…]
@@
The simplest construction of inversion sequences is the generating tree growing sequences on the right, used in e.g. [KMY24] or [Tes24, Section 2]. In comparison, growing sequences on the left is not the most practical way to construct the set of inversion sequences (it is not even obvious that the tree described by $\Omega_{\text{Left}}$ has $n!$ nodes at each level $n$). However, this construction becomes simpler for some classes of pattern-avoiding inversion sequences. For instance, [CMSW16] proves that the generating trees growing on the left for the classes $\mathcal{I}(101)$ and $\mathcal{I}(110)$ are both described by the succession rule
[…]
The generating tree growing on the left for inversion sequences lacks a property that is typically desired when working with objects that avoid patterns: it does not automatically restrict to a generating tree for families of pattern-avoiding inversion sequences. Specifically, restricting the tree to keep only sequences which avoid some patterns may disconnect the tree. This occurs because an inversion sequence which contains a pattern $\rho$ may have children which avoid $\rho$. For instance, the sequence $(0,0,2,1)$ avoids the pattern 010, but its parent sequence is $(0,1,0)$. We now characterize the patterns for which this restriction can be done, in Propositions 1 and 2. First, we show that this restriction is possible when avoiding a pattern which contains only one zero, or two zeros that are both at the beginning of the pattern.
@@
Let $\rho$ be a pattern of length $k\geqslant 1$ which either contains a single zero, or contains exactly two zeros and begins with 00. Let $\sigma$ be an inversion sequence which contains an occurrence $(\sigma_{q_{i}})_{i\in[1,k]}$ of $\rho$, and let $\tau$ be a child of $\sigma$ in the generating tree growing on the left. Then $\tau$ contains the pattern $\rho$.
@@
If $\rho$ contains a single zero, then $(\sigma_{q_{i}})_{i\in[1,k]}$ contains at most one zero. It follows that $(\sigma_{q_{i}})_{i\in[1,k]}$ is order-isomorphic to $(\tau_{q_{i}+1})_{i\in[1,k]}$ (every nonzero value is increased by 1, and the hypothetical zero can either increase by 1 or remain a zero), hence $\tau$ contains the pattern $\rho$.
If $\rho$ contains exactly two zeros and begins with 00, we distinguish some cases depending on the values of $(\sigma_{q_{i}})_{i\in[1,k]}$.
  • If $(\sigma_{q_{i}})_{i\in[1,k]}$ contains only positive values, then $(\tau_{q_{i}+1})_{i\in[1,k]}=(\sigma_{q_{i}}+1)_{i\in[1,k]}$ is still an occurrence of $\rho$, in $\tau$.
  • If $(\sigma_{q_{i}})_{i\in[1,k]}$ contains zero values, i.e. $\sigma_{q_{1}}=\sigma_{q_{2}}=0$, then $\tau_{q_{1}+1},\tau_{q_{2}+1}\in\{0,1\}$.
    • If $(\tau_{q_{1}+1},\tau_{q_{2}+1})\in\{(0,0),(1,1)\}$, then $(\tau_{q_{i}+1})_{i\in[1,k]}$ is still an occurrence of $\rho$.
    • If $(\tau_{q_{1}+1},\tau_{q_{2}+1})=(0,1)$, then $(\tau_{1})\cdot(\tau_{q_{i}+1})_{i\in\{1\}\cup[3,k]}$ is an occurrence of $\rho$.
    • If $(\tau_{q_{1}+1},\tau_{q_{2}+1})=(1,0)$, then $(\tau_{1})\cdot(\tau_{q_{i}+1})_{i\in[2,k]}$ is an occurrence of $\rho$. ∎
Proposition 1 provides sufficient conditions on a pattern $\rho$ to ensure that the complete generating tree growing on the left restricts to a generating tree for the class $\mathcal{I}(\rho)$. We now show that these conditions are also necessary.
Proposition 2.
Let $\rho$ be a pattern of length $k\geqslant 2$ which contains at least two zeros, and either begins with 000, or has a zero to the right of a positive value. Then there exist two inversion sequences $\sigma$ and $\tau$ such that
  • $\tau$ is a child of $\sigma$ in the generating tree growing on the left,
  • $\sigma$ contains $\rho$,
  • $\tau$ avoids $\rho$.
Proof.
If $\rho$ begins with 000, let $\rho^{\prime}=(\rho_{i}+k\cdot\chi(\rho_{i}>0))_{i\in[3,k]}$ be the sequence obtained by deleting the first two zeros of $\rho$ and adding $k$ to every nonzero value. Let $\sigma=(0,0,1,1,\dots,k-1,k-1)\cdot\rho^{\prime}$. It is easily checked that $\sigma$ is an inversion sequence which contains the pattern $\rho$. Let $\tau=\mathbf{child}(\sigma,\{1,2\})$. By construction, $\tau$ has one fewer zero than $\rho$, hence no occurrence of $\rho$ in $\tau$ can contain the value 0. Similarly, every value in $[1,k]$ appears exactly twice in $\tau$, while the value 0 appears at least three times in $\rho$, so any occurrence of $\tau$ in $\rho$ must only contain values greater than $k$. Only the rightmost k2k-2 entries of $\tau$ may have a value greater than $k$; this is not enough to form an occurrence of $\rho$.
If $\rho$ has a zero to the right of a positive value, let $\sigma=(0)^{k}\cdot\rho$. It is easy to see that $\sigma$ is an inversion sequence which contains $\rho$. Let $\tau=\mathbf{child}(\sigma,\{\mathbf{max}(\mathbf{Zeros}(\sigma))\})$. By hypothesis, the rightmost zero of $\sigma$ (or equivalently, $\rho)$ is necessarily an inversion bottom, i.e. there exists an entry of greater value to its left. This implies that the number of zeros that are inversion bottoms in $\tau$ is one fewer than in $\rho$, hence no occurrence of $\rho$ in $\tau$ can contain the value 0. Additionally, the value 1 appears only once in $\tau$, while the value 0 appears at least twice in $\rho$, so any occurrence of $\rho$ in $\tau$ must only contain values greater than 1. There are at most k2k-2 entries of $\tau$ having value at least 2; this is not enough to form an occurrence of $\rho$. ∎
By Proposition 1, the generating tree growing on the left is well-defined for inversion sequences avoiding the patterns 201 and 210 studied in Section 3, since both patterns contain a single zero. Even for problematic patterns like 010, the problem can be worked around by modifying our construction, in some cases. For instance, we describe a generating tree for the class $\mathcal{I}(010,102)$ in Section 4. Some other workarounds for the classes $\mathcal{I}(000)$, $\mathcal{I}(010)$, and $\mathcal{I}(100)$ can be found in Section 5.
@@ 3 Inversion sequences avoiding 201 and 210
[…]
@@
The proposition above establishes that in the generating tree growing on the left for $\mathcal{I}(201,210)$, the children of $\sigma$ are the sequences such that the subset $Z$ of zeros to which the value 1 is added must contain either all or none [^3] When $R(\sigma)$ is not empty, one can think of it as containing only a single zero, since all zeros in this set are modified in the same way. of $R(\sigma)$. Any subset of $L(\sigma)$ may intersect with $Z$ without creating an occurrence of 201 or 210. It follows that the number of children of $\sigma$ is $2^{\ell(\sigma)+\chi(R(\sigma)\neq\emptyset)}$.
@@
    • If $Z$ is a (possibly empty) interval of positions at the end of $L(\sigma)$, that is if $Z=[i,\ell(\sigma)]$ for some $i\in[1,\ell(\sigma)+1]$ (the case $i=\ell(\sigma)+1$ corresponding to $Z=\emptyset$), then $L(\tau)=[1,i]$ and $R(\tau)=\emptyset$, therefore $s(\tau)=i$.
[…]
@@
    • Otherwise, $Z$ is not empty nor consists of a set of final zeros of $\sigma$. This implies as before that $R(\tau)\neq\emptyset$, and that $s(\tau)=i+1$, where $i=\ell(\tau)=\mathbf{min}(Z)\in[1,\ell(\sigma)]$. There are $2^{s(\sigma)-i}-1$ possible sets $Z$ satisfying the present assumptions: either $Z=\{i\}\sqcup Z^{\prime}$ for some $Z^{\prime}\subseteq[i+1,\ell(\sigma)]$, or $Z=\{i\}\sqcup Z^{\prime}\sqcup R(\sigma)$ for some $Z^{\prime}\subsetneq[i+1,\ell(\sigma)]$). ∎
@@
We apply the kernel method. The kernel in (1) is $1+\frac{tu^{2}}{2-u}$. To cancel the kernel, we substitute $u$ for $U(t)=\frac{1-\sqrt{1-8t}}{2t}$. This yields the equation